OpenStudy (anonymous):

if y=1/2 and x=-14, find y.

OpenStudy (lgbasallote):

what's the relationship between x and y

OpenStudy (anonymous):

y=1/2 x

OpenStudy (lgbasallote):

substitute then

OpenStudy (lgbasallote):

x = -14 therefore y = 1/2 (-14)

OpenStudy (anonymous):

1/2 . (-14) = y

OpenStudy (lgbasallote):

that's also equal to \[-\frac{14}{2} = -14 \div 2\]

OpenStudy (anonymous):

so it will be -7

OpenStudy (lgbasallote):

yep